# Efficient Policy Optimization in Robust Constrained MDPs with Iteration Complexity Guarantees

### v2

@@ Abstract
Constrained decision-making is essential for designing safe policies in real-world control systems, yet simulated environments often fail to capture real-world adversities. We consider the problem of learning a policy that will maximize the cumulative reward while satisfying a constraint, even when there is a mismatch between the real model and an accessible simulator/nominal model. In particular, we consider the robust constrained Markov decision problem (RCMDP) where an agent needs to maximize the reward and satisfy the constraint against the worst possible stochastic model under the uncertainty set centered around an unknown nominal model. Primal-dual methods, effective for standard constrained MDP (CMDP), are not applicable here because of the lack of the strong duality property. Further, one cannot apply the standard robust value-iteration based approach on the composite value function either as the worst case models may be different for the reward value function and the constraint value function. We propose a novel technique that effectively minimizes the constraint value function–to satisfy the constraints; on the other hand, when all the constraints are satisfied, it can simply maximize the robust reward value function. We prove that such an algorithm finds a policy with at most $\epsilon$ sub-optimality and feasible policy after $O(\epsilon^{-2})$ iterations. In contrast to the state-of-the-art method, we do not need to employ a binary search, thus, we reduce the computation time by at least 4x for smaller value of discount factor ($\gamma$) and by at least 6x for larger value of $\gamma$.
@@ 1 Introduction
Hence, the objective is passed on to the constraint with an objective of how tight the constraint can be. [8] finds the optimal policy for each $b_{0}$, and then optimized $b_{0}$ using a binary search. They showed that for each $b_{0}$, the iteration complexity is $O(\epsilon^{-4})$ to find the optimal policy. Note that one needs to evaluate robust value function at every iteration for each $b_{0}$ which is costly operation especially when $\gamma$ is large as it is evident by Table 1. Further, the binary search method only works when the estimation is perfect [9], thus, if the robust policy evaluator is noisy which is more likely for the large state-space, the binary search method may not work. Moreover, the complexity of iteration is only $O(\log(\epsilon^{-1})\epsilon^{-4})$, which is worse than that of the CMDP [10]. We seek to answer the following:
@@
CRS 48.574 190.53 228.65 301.21
Garnet 78.406 390.21 516.23 653.14
Table 1: Comparison of execution times averaged over multiple runs between RNPG, and EPIRC-PGS (inner loop T = 100 and outer-loop K=10)
@@
  • Our empirical results show that our proposed approaches achieve a feasible policy with good reward (comparable or better than the one achieved by EPIRC-PGS) at a faster wall-clock time (see Table 1) [^1] The system specifications are, Processor: Intel(R)Core(TM)i7-14700-2.10 GHz, Installed RAM 32.0 GB (31.7 GB usable),64-bit operating system, x64-based processor No GPU. compared to the EPIRC-PGS. From Table 1, it is evident that our algorithm speeds up the computation process by at-least 4 times as compared to EPIRC-PGS algorithm when $\gamma=0.9$ and at-least 6 times to EPIRC-PGS when $\gamma=0.995$.
@@ 1.1 Related Works
CMDP: The convex nature of the state-action occupancy measure ensures the existence of a zero duality gap between the primal and dual problem for CMDP, making them well-suited for solution via primal-dual methods [2, 11, 12, 13, 14, 15, 16, 17, 18]. The convergence bounds and rates of convergence for these methods have been extensively studied in [19, 20, 21, 22, 23, 6]. Beyond primal-dual methods, LP-based and model-based approaches have been explored to solve the primal problem directly [24, 17, 25, 26, 27, 28]. However, the above approaches cannot be extended to the RCMDP case.
@@ 2 Problem Formulation
In the CMDP problem, [44] shows that the strong duality holds when there exists a strictly feasible policy (aka Slater’s condition). However, [7] highlighted that strong duality may not hold for the RCMDP problem as the occupancy measure is no longer convex as the worst transition model differs for different policies. In addition to that, in [8] a strong ambiguity regarding the tractability in solving lagrangian problem is discussed. Further, even if one fixes $\lambda$, one cannot apply robust value iteration approach to find the optimal policy for the Lagrangian unlike the CMDP. Hence, it is evident to look for alternative measures to find a solution to the optimality problem.
@@ 3.1 Our Proposed Approach
Intuition: Note that when $J_{c_{i}}^{\pi}\leq b_{i}$ for all $i=1,\ldots,K$, the second term in the objective becomes negative, and since $J_{c_{0}}^{\pi}\geq 0$, the optimization will focus on minimizing $J_{c_{0}}^{\pi}$, as the policy is likely to be feasible with respect to all constraints.
Conversely, if there exists any $i$ such that $J_{c_{i}}^{\pi}>b_{i}$, then for a sufficiently large $\lambda$, the term $J_{c_{0}}^{\pi}/\lambda$ becomes smaller than $J_{c_{i}}^{\pi}-b_{i}$, causing the optimization to prioritize reducing the most violated constraint $J_{c_{i}}^{\pi}$. Even though we can not claim that (9) and (1) are the same, we can claim that the optimal solution of (9) can only violate the constraint by at most $\epsilon$-amount by a suitable choice of $\lambda$. Hence, minimizing (9) amounts to searching for policies that can violate at most $\epsilon$ amount. Thus, the optimal policy of (1) can be an optimal of (9). In particular, optimal policy of (9) indeed has a smaller cost compared to that of (1). We formalize this as the following result.
@@
Note that even though for theoretical analysis to achieve a feasible policy we assume the knowledge of $\xi$, for our empirical evaluations, we did not assume that and yet we achieved feasible policy with good reward exceeding the state-of-the-art performance. Hence, we modify the policy space to be $\xi$-dependent.
@@ 3.2 Policy Optimization Algorithm
The robust value function under the KL-uncertainty set is formalized in Lemma 38 (see Appendix D). The advantage is that we obtain a closed form expression for the robust value function, ans we can evaluate it by drawing samples from the nominal model only. For further background on KL and other uncertainty sets, we refer the reader to [47, 32, 31].
@@
Consider $i_{t}=\arg\min\{J_{c_{0}}^{\pi_{t}}/\lambda,\{J_{c_{i}}^{\pi_{t}}-b_{i}+\xi\}_{i=1}^{K}\}$, and $p_{t}=\arg\max J_{c_{i_{t}}}^{\pi_{t},p_{t}}$, we can evaluate $\nabla_{\pi_{t}}J_{c_{i_{t}}}^{\pi_{t},p_{t}}$ using the robust evaluator for $Q_{c_{i_{t}}}^{\pi_{t},p_{t}}(\cdot,\cdot)$ as mentioned.
@@
Output policy $\arg\min_{t\in 0,\ldots,T-1}\max\{J_{c_{0}}^{\pi_{t}}/\lambda,\max_{i}(J_{c_{i}}^{\pi_{t}}-b_{i}+\xi)\}$.
@@ 4 Theoretical Results
This was a common assumption for unconstrained RMDP as well [50, 51]. More importantly, Algorithm 1 does not need to know $\beta$.
@@
Under Assumptions 1 and 2, with $\lambda=2H/\min\{\xi,1\}$, $\alpha_{t}=\alpha_{0}=\dfrac{1-\gamma}{\sqrt{TS}}$ after $T=O(\epsilon^{-2}\xi^{-2}(1-\gamma)^{-2}(1-\beta)^{-2}S\log(|A|))$ iterations, Algorithm 1 returns a policy $\hat{\pi}$ such that $J_{r}^{\hat{\pi}}-J_{r}^{\pi^{*}}\leq\epsilon$, and $\max_{n}J_{c_{n}}^{\hat{\pi}}-b_{n}\leq 0$.
@@ 4.1 Proof Outline
for any policy $\pi$ after $O(\lambda^{2}\epsilon^{-2}(1-\gamma)^{-4}(1-\beta)^{-2})$ iterations under Assumption 2.
@@ 5 Experimental Results
We evaluate our algorithms [^2] The complete implementation is available at https://github.com/Sourav1429/RCAC_NPG on two environments: (i) Garnet, and (ii) Constrained Riverswim (CRS). Additional experimental results are provided in Appendix E.
@@ 5.1 Analysis of results
    Computational Time: RNPG exhibits significant improvements in computational efficiency, achieving convergence at least 4x faster than EPIRC-PGS for $\gamma=0.9$, and 6x faster for $\gamma=0.995$ in the CRS setting (Table 1). In the Garnet environment, RNPG achieves a 5x speedup over EPIRC-PGS for $\gamma=0.9$, and at least 8x speedup for $\gamma=0.995$ (Table 1). The higher speed up in the Garnet environment because of the higher state-space. The difference in runtime can be attributed to the fact that RNPG eliminates the need for binary search for each $b_{0}$ value in (2) as described above, and it uses a KL regularization.
@@
    We also compare RNPG with RPPG (see Appendix D), a projected robust gradient descent variant that uses an $\ell_{2}$ regularizer instead of KL for policy update in (11). In the CRS environment, RPPG performs slightly better than EPIRC-PGS by maintaining smaller constraint violations, though it still occasionally breaches the safety threshold. In the Garnet environment, RPPG achieves a better performance compared to EPIRC-PGS while maintaining feasibility, however, it achieves a smaller reward compared to RNPG. RNPG is also much stable, showing that KL regularization is more effective compared to $\ell_{2}$ regularization. Note that RPPG also has a smaller computational time compared to EPIRC-PGS (see Table 11), which demonstrates that removing the binary search is the key, as EPIRC-PGS also uses $\ell_{2}$ regularization for policy update.
@@ 6 Discussions and Limitation
Limitation: A major limitation of our work is that we assumed finite state-space. The extension to infinite state-space constitutes an important future research direction. Other limitation is that Assumption 2. Relaxing Assumption 2 constitutes an important future research direction. Achieving a lower bound or improving the iteration complexity is also an important future research direction.
7 Conclusions
@@ Acknowledgments
Kishan Panaganti and Adam Wierman acknowledge support from the Resnick Institute and the ‘PIMCO Postdoctoral Fellow in Data Science’ fellowship at Caltech. Arnob Ghosh and Sourav Ganguly acknowledge support from the NJIT faculty startup fund.
@@ Appendix A Proof of Proposition 1
$\Phi(\pi)-\Phi(\pi^{*})\leq\dfrac{1}{1-\beta}\mathbb{E}_{s\sim d_{\rho}^{\pi^{*},p_{0}}}[\langle Q^{\pi,p}_{c_{i}},\pi_{s}-\pi^{*}],$ (20)
@@
The iterate update pair $(i_{t},p_{t})$ satisfies the following:
$\Phi(\pi_{t})-\Phi(\pi^{*})=\max\{J_{c_{0}}^{\pi_{t}}-b_{0},J_{c_{i}}^{\pi_{t}}-b_{i}\}-\max\{J_{c_{0}}^{\pi^{*}}-b_{0},J_{c_{i}}^{\pi^{*}}-b_{i}\}$
$\leq J_{c_{i_{t}}}^{\pi_{t}}-J_{c_{i_{t}}}^{\pi^{*}}\leq J_{c_{i_{t}}}^{\pi_{t},p_{t}}-J_{c_{i_{t}}}^{\pi^{*},p_{t}}.$ (21)
We assume that the robust policy evaluator is $\epsilon_{0}$-accurate, hence, we have from (B)
$\Phi(\pi_{t})-\Phi(\pi^{*})=J_{c_{i_{t}}}^{\pi_{t},p_{t}}-J_{c_{i_{t}}}^{\pi^{*},p_{t}}$
$\leq J_{c_{i_{t}}}^{\pi_{t},p_{t}}-J_{c_{i_{t}}}^{\pi^{*},p_{t}}+\epsilon_{0}.$ (22)
Here, we use the following result from Lemma B.3
$J_{c_{i_{t}}}^{\pi_{t},p_{t}}-J_{c_{i_{t}}}^{\pi^{*},p_{t}}\leq\dfrac{1}{1-\beta}\sum_{s}d_{\rho}^{\pi^{*},p_{0}}(s)\sum_{a}(\pi_{t,s,a}-\pi^{*}_{s,a})Q_{s,a,c_{i_{t}}}^{\pi_{t},p_{t}}.$
From Lemma B.1, we then have from (B)
$\Phi(\pi_{t})-\Phi(\pi^{*})\leq$ (23)
$[\dfrac{1}{1-\beta}\mathbb{E}_{s\sim d_{\rho}^{\pi^{*},p_{0}}}[\langle Q_{s,i_{t}}^{\pi_{t},p_{t}},\pi_{t,s}-\pi_{t+1,s}\rangle+\dfrac{1}{\alpha}B(\pi^{*},\pi_{t})-\dfrac{1}{\alpha}B(\pi^{*},\pi_{t+1})-\dfrac{1}{\alpha}B(\pi_{t+1},\pi_{t})]]+\epsilon_{0}.$ (24)
@@
Hence, by summing over $t$, and using (B) we have from (23),
@@
$+\dfrac{1}{\alpha(1-\beta)}\mathbb{E}_{s\sim d_{\rho}^{\pi^{*},p_{0}}}[B(\pi^{*},\pi_{t})-B(\pi^{*},\pi_{t+1})]+T\epsilon_{0}$
$\leq\dfrac{1}{(1-\beta)}\sum_{t=0}^{T-1}\alpha S\dfrac{1}{(1-\gamma)^{2}}+\dfrac{1}{\alpha(1-\beta)}\mathbb{E}_{s\sim d_{\rho}^{\pi^{*},p_{0}}}B(\pi^{*},\pi_{0})+T\epsilon_{0}.$ (26)
@@
$\sum_{t}(\Phi(\pi_{t})-\Phi(\pi^{*}))\leq\dfrac{1}{(1-\beta)}T\alpha S\dfrac{1}{(1-\gamma)^{2}}+\dfrac{1}{\alpha(1-\beta)}\mathbb{E}_{s\sim d_{\rho}^{\pi^{*},p_{0}}}B(\pi^{*},\pi_{0})+T\epsilon_{0}.$ (27)
@@
$\sum_{t}(\Phi(\pi_{t})-\Phi(\pi^{*}))\leq\dfrac{1}{1-\beta}\sqrt{ST}\log(|A|)\dfrac{1}{\sqrt{(1-\gamma)^{2}}}+T\epsilon_{0}.$ (28)
@@
$\Phi(\hat{\pi})-\Phi(\pi^{*})\leq\dfrac{1}{T}\sum_{t}(\Phi(\pi_{t})-\Phi(\pi^{*}))\leq\dfrac{\sqrt{S}\log(|A|)}{(1-\beta)\sqrt{T(1-\gamma)^{2}}}+\epsilon_{0},$ (29)

### v3

@@ Abstract
Constrained decision-making is essential for designing safe policies in real-world control systems, yet simulated environments often fail to capture real-world adversities. We consider the problem of learning a policy that will maximize the cumulative reward while satisfying a constraint, even when there is a mismatch between the real model and an accessible simulator/nominal model. In particular, we consider the robust constrained Markov decision problem (RCMDP) where an agent needs to maximize the reward and satisfy the constraint against the worst possible stochastic model under the uncertainty set centered around an unknown nominal model. Primal-dual methods, effective for standard constrained MDP (CMDP), are not applicable here because of the lack of the strong duality property. Further, one cannot apply the standard robust value-iteration based approach on the composite value function either as the worst case models may be different for the reward value function and the constraint value function. We propose a novel technique that effectively minimizes the constraint value function–to satisfy the constraints; on the other hand, when all the constraints are satisfied, it can simply maximize the robust reward value function. We prove that such an algorithm finds a policy with at most $\epsilon$ sub-optimality and feasible policy after $O(\epsilon^{-2})$ iterations. In contrast to the state-of-the-art methods, we do not need to employ a binary search, thus, we reduce the computation time for larger value of discount factor ($\gamma$), and achieve a better performance for large state space.
@@ 1 Introduction
Hence, the objective is passed on to the constraint with an objective of how tight the constraint can be. [8] finds the optimal policy for each $b_{0}$, and then optimized $b_{0}$ using a binary search. They showed that for each $b_{0}$, the iteration complexity is $O(\epsilon^{-4})$ to find the optimal policy. Note that one needs to evaluate robust value function at every iteration for each $b_{0}$ which is costly operation especially when $\gamma$ is large as it is evident by Table 1. Further, the binary search method only works when the estimation is perfect [9], thus, if the robust policy evaluator is noisy which is more likely for the large state-space, the binary search method may not work as it is evident in our function approximation setup (Appendix G). Moreover, the complexity of iteration is only $O(\log(\epsilon^{-1})\epsilon^{-4})$, which is worse than that of the CMDP [10]. We seek to answer the following:
@@
CRS 48.574 190.53 228.65 290.15
Garnet 78.406 290.21 316.23 453.14
Modified Frozenlake 160.31 453.7 561.47 620.12
Garbage collector 177.13 344.87 400.13 489.54
Table 1: Comparison of execution times averaged over multiple runs between RNPG, and EPIRC-PGS (inner loop T = 100 and outer-loop K=10) (Some more experimental results to demonstrate faster performance of our algorithms can be found in appendix G)
@@
  • We extend our framework to the function approximation setup by proposing a robust constrained actor-critic with integral probability metric as the uncertainty metric. For the finite-state, our empirical results show that our proposed approaches achieve a feasible policy with good reward (comparable or better than the one achieved by EPIRC-PGS, see Table 11) at a faster wall-clock time (see Table 1) [^1] The system specifications are, Processor: Intel(R)Core(TM)i7-14700-2.10 GHz, Installed RAM 32.0 GB (31.7 GB usable),64-bit operating system, x64-based processor No GPU. compared to the EPIRC-PGS. From Table 1, it is evident that our algorithm speeds up the computation process by at-least 2 times as compared to EPIRC-PGS algorithm when $\gamma=0.9$ and at-least 3 times to EPIRC-PGS when $\gamma=0.995$. For the function approximation setup, our proposed approach is the only one that achieves feasibility and even a better reward to the robust version of CRPO [11] during the test time for Cartpole experiment (Table 12). Further, we outperform EPIRC-PGS significant manner for the function approximation setup both in terms of performance and the training time showing its efficacy.
@@ 1.1 Related Works
CMDP: The convex nature of the state-action occupancy measure ensures the existence of a zero duality gap between the primal and dual problem for CMDP, making them well-suited for solution via primal-dual methods [2, 12, 13, 14, 15, 16, 17, 18, 19]. The convergence bounds and rates of convergence for these methods have been extensively studied in [20, 21, 22, 23, 24, 6, 25, 26]. Beyond primal-dual methods, LP-based and model-based approaches have been explored to solve the primal problem directly [27, 18, 28, 29, 11, 30]. However, the above approaches cannot be extended to the RCMDP case.
@@ 2 Problem Formulation
In the CMDP problem, [46] shows that the strong duality holds when there exists a strictly feasible policy (aka Slater’s condition). However, a concurrent work [47] highlighted that strong duality does not hold for the RCMDP problem as the occupancy measure is no longer convex as the worst transition model differs for different policies. In addition to that, in [8] a strong ambiguity regarding the tractability in solving lagrangian problem is discussed. Further, even if one fixes $\lambda$, one cannot apply robust value iteration approach to find the optimal policy for the Lagrangian unlike the CMDP. Hence, it is evident to look for alternative measures to find a solution to the optimality problem.
@@ 3.1 Our Proposed Approach
Intuition: Note that when $J_{c_{i}}^{\pi}\leq b_{i}$ for all $i=1,\ldots,K$, the second term in the objective becomes negative, and since $J_{c_{0}}^{\pi}\geq 0$, the optimization will focus on minimizing $J_{c_{0}}^{\pi}$, as the policy is likely to be feasible with respect to all constraints. Conversely, if there exists any $i$ such that $J_{c_{i}}^{\pi}>b_{i}$, then for a sufficiently large $\lambda$, the term $J_{c_{0}}^{\pi}/\lambda$ becomes smaller than $J_{c_{i}}^{\pi}-b_{i}$, causing the optimization to prioritize reducing the most violated constraint $J_{c_{i}}^{\pi}$.
Even though we can not claim that (9) and (1) are the same, we can claim that the optimal solution of (9) can only violate the constraint by at most $\epsilon$-amount by a suitable choice of $\lambda$. Hence, minimizing (9) amounts to searching for policies that can violate at most $\epsilon$ amount. Thus, the optimal policy of (1) can be an optimal of (9). In particular, optimal policy of (9) indeed has a smaller cost compared to that of (1). We formalize this as the following result.
@@
Note that even though for theoretical analysis to achieve a feasible policy we assume the knowledge of $\xi$; for our empirical evaluations, we did not assume that and yet we achieved feasible policy with good reward exceeding the state-of-the-art performance. Hence, we modify the policy space to be $\xi$-dependent.
@@ 3.2 Policy Optimization Algorithm
The robust value function under the KL-uncertainty set is formalized in Lemma 37 (see Appendix D). The advantage is that we obtain a closed form expression for the robust value function, and we can evaluate it by drawing samples from the nominal model only. For further background on KL and other uncertainty sets, we refer the reader to [50, 34, 33].
@@
Consider $i_{t}=\arg\max\{J_{c_{0}}^{\pi_{t}}/\lambda,\{J_{c_{i}}^{\pi_{t}}-b_{i}+\xi\}_{i=1}^{K}\}$, and $p_{t}=\arg\max J_{c_{i_{t}}}^{\pi_{t},p_{t}}$, we can evaluate $\nabla_{\pi_{t}}J_{c_{i_{t}}}^{\pi_{t},p_{t}}$ using the robust evaluator for $Q_{c_{i_{t}}}^{\pi_{t},p_{t}}(\cdot,\cdot)$ as mentioned.
@@
Output policy $\arg\min_{\pi_{t}s.t.t\in\{0,\ldots,T-1\}}\max\{J_{c_{0}}^{\pi_{t}}/\lambda,\max_{i}(J_{c_{i}}^{\pi_{t}}-b_{i}+\xi)\}$.
@@ 4 Theoretical Results
This was a common assumption for unconstrained RMDP as well [53, 54]. Assumption 2 states that if the perturbed distribution assigns positive probability to an event, the nominal model should also assign positive probability to that event. Otherwise, a mismatch in supports could lead to unsampled regions and render finite‑iteration bounds intractable. More importantly, Algorithm 1 does not need to know $\beta$. We also did not enforce in our empirical studies. The algorithm still performed well, suggesting that the practical impact may be less restrictive than the theory implies. Also, EPIRC-PGS [8] assumed that the ratio between the state-action occupancy measures on the states covered by all policies and the initial state distribution is bounded.
We also consider a slightly stronger optimal policy for the surrogate problem.
Assumption 3.
We consider $\hat{\pi}^{*}$,a uniform minimizer across all states of the surrogate problem in (9), i.e., $\hat{\pi}^{*}$ is a solution of $\min_{\pi}\max\{V_{c_{0}}^{\pi,P}(s)/\lambda,\max_{n}\max_{P}V_{c_{i}}^{\pi,P}-b_{n}\}$ for all $s$.
A similar assumption is also considered for the unconstrained problem [54, 31].
@@
Under Assumptions 1, 2, and 3 with $\lambda=2H/\min\{\xi,1\}$, $\alpha_{t}=\alpha_{0}=\dfrac{1-\gamma}{\sqrt{TS}}$ after $T=O(\epsilon^{-2}\xi^{-2}(1-\gamma)^{-2}(1-\beta)^{-2}S\log(|A|))$ iterations, Algorithm 1 returns a policy $\hat{\pi}$ such that $J_{r}^{\hat{\pi}}-J_{r}^{\pi^{*}}\leq\epsilon$, and $\max_{n}J_{c_{n}}^{\hat{\pi}}-b_{n}\leq 0$.
@@ 4.1 Proof Outline
for any policy $\pi$ after $O(\lambda^{2}\epsilon^{-2}(1-\gamma)^{-4}(1-\beta)^{-2})$ iterations under Assumptions 2, and 3.
@@ 5 Experimental Results
We evaluate our algorithms [^2] The complete implementation is available at https://github.com/Sourav1429/RCAC_NPG.git on two environments: (i) Garnet, and (ii) Constrained Riverswim (CRS). Additional experimental results are provided in Appendix E.
@@ 5.1 Analysis of results
    Computational Time: RNPG exhibits significant improvements in computational efficiency, achieving convergence at least 4x faster than EPIRC-PGS for $\gamma=0.9$, and 6x faster for $\gamma=0.995$ in the CRS setting (Table 1). In the Garnet environment, RNPG achieves a 3x speedup over EPIRC-PGS for $\gamma=0.9$, and at least 5x speedup for $\gamma=0.995$ (Table 1). The difference in runtime can be attributed to the fact that RNPG eliminates the need for binary search for each $b_{0}$ value in (2) as described above, and it uses a KL regularization.
@@
    We also compare RNPG with RPPG (see Appendix D), a projected robust gradient descent variant that uses an $\ell_{2}$ regularizer instead of KL for policy update in (11). In the CRS environment, RPPG performs slightly better than EPIRC-PGS by maintaining smaller constraint violations, though it still occasionally breaches the safety threshold. In the Garnet environment, RPPG achieves a better performance compared to EPIRC-PGS while maintaining feasibility, however, it achieves a smaller reward compared to RNPG. RNPG is also much stable, showing that KL regularization is more effective compared to $\ell_{2}$ regularization. We observe that RPPG (and, similar to RNPG) also has a smaller computational time compared to EPIRC-PGS, which demonstrates that removing the binary search is the key, as EPIRC-PGS also uses $\ell_{2}$ regularization for policy update.
@@ 6 Discussions and Limitation
6.1 Extending to Function Approximation: Robust Constrained Actor-Critic (RCAC)
We extend our framework to the function approximation setting motivated by the work of [54] for unconstrained Robust MDP problem. In particular, we consider the integral probability metric (IPM) as an uncertainty set, $d_{\mathcal{F}}(p,q)=\sup_{f\in\mathcal{F}}\{p^{T}f-q^{T}f\}$ where $\mathcal{F}\subseteq\mathbb{R}^{|S|}$. Many metrics such as Kantorovich metric, total variation, etc., are special cases of IPM under different function classes [56]. We then consider the IPM-based uncertainty set, $\mathcal{P}_{s,a}=\{q|d_{\mathcal{F}}(q,p^{0}_{s,a})\leq\rho\}$ around the nominal model.
[…]
with the next state $s^{\prime}$ is drawn from the nominal model. Guided by the last regularization term of the empirical robust Bellman operator in (15), when considering value function approximation by neural networks we add a similar regularization term for all the neural network parameters except for the bias parameter in the last layer. We use the expression in (15) for the robust value function for the gradient, and $J_{c_{i}}$ in Algorithm 1. We need to estimate the robust Q function. In order to estimate the $Q$-function we first use (15) by plugging the $V$-approximation, and then we use the linear regression to fit the critic for the robust $Q$-function. The details can be found in Appendix G.
From the empirical results in Figure 9 and Table 12, it is evident that our proposed approach outperforms the state-of-the-art approaches. More importantly, compared to the EPIRC-PGS (adapted to the function approximation setting), our approach achieves a significantly better performance with small wall-clock time. Also, our proposed approach outperforms the robust version of the CRPO algorithm [11, 47] and achieves feasibility unlike robust CRPO. In Appendix H we showed that robust CRPO may not achieve a finite time iteration complexity guarantee even for finite-state space.
7 Conclusions and Future Works
@@
Relaxing Assumption 2, and 3 constitute an important future research direction. Achieving a lower bound or improving the iteration complexity is also an important future research direction. Characterizing the results for other uncertainty sets also constitutes an important future research direction. Iteration complexity guarantee for the function approximation setting has been left for the future.
@@ Acknowledgments
AG and SG acknowledge NJIT Startup Fund indexed 172884. SG acknowledges Neurips 2025 for awarding him with the NeurIPS 2025 Scholar Award. AW and KP acknowledge NSF grants CCF-2326609, CNS-2146814, CPS-2136197, CNS-2106403, and NGSDI-2105648 and support from the Resnick Sustainability Institute. KP also acknowledges support from the ‘PIMCO Postdoctoral Fellow in Data Science’ fellowship at the California Institute of Technology and the Resnick Institute.
@@ Appendix A Proof of Proposition 1
For any $\pi\in\Pi$
$\Phi(\pi)-\Phi(\hat{\pi}^{*})\leq\dfrac{1}{1-\beta}\mathbb{E}_{s\sim d_{\rho}^{\pi^{*},p_{0}}}[\langle Q^{\pi,p}_{c_{i}},\pi_{s}-\hat{\pi}^{*}\rangle],$ (21)
@@
From Lemma B.3
$\Phi(\pi_{t})-\Phi(\hat{\pi}^{*})\leq\dfrac{1}{1-\beta}\sum_{s}d_{\rho}^{\hat{\pi}^{*},p_{0}}(s)\sum_{a}(\pi_{t,s,a}-\hat{\pi}^{*}_{s,a})\hat{Q}_{s,a,c_{i_{t}}}^{\pi_{t},p_{t}}.$
where $\hat{Q}$ is the estimated value. Consider that the worst-case evaluator is only $\epsilon_{0}$ is close that is $Q_{c_{i_{t}}}^{\pi_{t},p_{t}}-\hat{Q}_{c_{i_{t}}}^{\pi_{t},p_{t}}||_{\infty}\leq\epsilon_{0}$.
Hence, we have
$\Phi(\pi_{t})-\Phi(\hat{\pi}^{*})\leq\dfrac{1}{1-\beta}\sum_{s}d_{\rho}^{\hat{\pi}^{*},p_{0}}(s)\sum_{a}(\pi_{t,s,a}-\hat{\pi}^{*}_{s,a})Q_{s,a,c_{i_{t}}}^{\pi_{t},p_{t}}+\epsilon_{0}$ (22)
Applying Lemma B.1 (subtracting and adding $\langle Q_{s,i_{t}}^{\pi_{t},p_{t}},\pi_{t,s}\rangle$ , we then have from (22)
$\Phi(\pi_{t})-\Phi(\hat{\pi}^{*})\leq$
$[\dfrac{1}{1-\beta}\mathbb{E}_{s\sim d_{\rho}^{\hat{\pi}^{*},p_{0}}}[\langle Q_{s,i_{t}}^{\pi_{t},p_{t}},\pi_{t,s}-\pi_{t+1,s}\rangle+\dfrac{1}{\alpha}B(\hat{\pi}^{*},\pi_{t})-\dfrac{1}{\alpha}B(\hat{\pi}^{*},\pi_{t+1})-\dfrac{1}{\alpha}B(\pi_{t+1},\pi_{t})]]+\epsilon_{0}.$ (23)
@@
Hence, by summing over $t$, and using (B) we have from (B),
@@
$+\dfrac{1}{\alpha(1-\beta)}\mathbb{E}_{s\sim d_{\rho}^{\hat{\pi}^{*},p_{0}}}[B(\hat{\pi}^{*},\pi_{t})-B(\hat{\pi}^{*},\pi_{t+1})]+T\epsilon_{0}$
$\leq\dfrac{1}{(1-\beta)}\sum_{t=0}^{T-1}\alpha S\dfrac{1}{(1-\gamma)^{2}}+\dfrac{1}{\alpha(1-\beta)}\mathbb{E}_{s\sim d_{\rho}^{\hat{\pi}^{*},p_{0}}}B(\hat{\pi}^{*},\pi_{0})+T\epsilon_{0}.$ (25)
@@
$\sum_{t}(\Phi(\pi_{t})-\Phi(\hat{\pi}^{*}))\leq\dfrac{1}{(1-\beta)}T\alpha S\dfrac{1}{(1-\gamma)^{2}}+\dfrac{1}{\alpha(1-\beta)}\mathbb{E}_{s\sim d_{\rho}^{\hat{\pi}^{*},p_{0}}}B(\hat{\pi}^{*},\pi_{0})+T\epsilon_{0}.$ (26)
@@
$\sum_{t}(\Phi(\pi_{t})-\Phi(\hat{\pi}^{*}))\leq\dfrac{1}{1-\beta}\sqrt{ST}\log(|A|)\dfrac{1}{\sqrt{(1-\gamma)^{2}}}+T\epsilon_{0}.$ (27)
@@
$\Phi(\hat{\pi})-\Phi(\hat{\pi}^{*})\leq\dfrac{1}{T}\sum_{t}(\Phi(\pi_{t})-\Phi(\hat{\pi}^{*}))\leq\dfrac{\sqrt{S}\log(|A|)}{(1-\beta)\sqrt{T(1-\gamma)^{2}}}+\epsilon_{0},$ (28)